OpenStudy (anonymous):

Integrate sin(x^2) wrt x.

OpenStudy (shubhamsrg):

hint: use this cos2x = 1- 2sin^2x ;)

OpenStudy (anonymous):

do u know that sin^2a=(1-cos2a)/2

OpenStudy (anonymous):

isn't that\[\int \sin x^2 dx\]?

OpenStudy (shubhamsrg):

ohh..hahaha..lol.. really sorry...must be what you are saying..hmm..

OpenStudy (anonymous):

there is no closed form for this integral...we can do it numerically...

OpenStudy (anonymous):

substitute sinx^2 with (1-cos2x)/2

OpenStudy (shubhamsrg):

yes,,no elementary soln..

OpenStudy (shubhamsrg):

well if you apply limits surely an ans is there..

OpenStudy (anonymous):

but there is an interval that this integral can be done analytically...and that is\[\int_{0}^{\infty} \sin x^2 dx\]

OpenStudy (anonymous):

o.0 just substitute sinx^2 with (1-cos2x)/2 and integrate two parts ... (1/2)dx and (cos2x/2)dx

OpenStudy (anonymous):

@protosbg thats \(\sin x^2\) not \(\sin^2 x\)

OpenStudy (anonymous):

lol :D sorry

OpenStudy (shubhamsrg):

how about simpson's rule here?? we may take limits a to b,,any arbitrary constants..

OpenStudy (anonymous):

Mukushla, what is the proof of the thing after <<we know that>>?

OpenStudy (anonymous):

we must make it a double integral...and change it from cartesian to polar...

OpenStudy (anonymous):

and we will treat \(t\) like a constant here

OpenStudy (anonymous):